OpenStudy (anonymous):

How do I simplify the equation and write it in standard form. 192 = x (x-4) (x-10)

OpenStudy (anonymous):

192=x(x^2-14x+40) 192=x^3-14x^2+40x 0=x^3-14x^2+40x-192

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem